OpenStudy (anonymous):

Can someone help m do this. 4(15-5r) =90

OpenStudy (gk99):

60 - 20r = 90 -20r = 30 r = -3/2

OpenStudy (gunnar):

ok add 15 to 5r it comes out as 20 so then you divide 20/4= 5 so then divide 90 by 5 90/5=23